# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ 4. Maxwellian bound for the cut-off case for non-fully specular reflection condition
Notice that we used the fact that $\lambda\leq 1$. Hence, by defining $\mathcal{A}_{2}(\tau_{2}):=\lambda^{2}\mathcal{A}_{1}(\tau_{2})$, we conclude the proof of Theorem 1.1 in the case $0\leq\alpha<1$.
@@ 5. Maxwellian bound in the cut-off case for fully specular reflection condition
In this chapter, we consider the fully specular reflection condition $\alpha=1$. Throughout the chapter, some of the proof is omitted because of the similarity as in Chapters 3 and 4. For a detailed proof, see .
@@
(5.3) $\begin{cases}a^{S}_{0}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=a_{0}(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\langle\mathfrak{R}\rangle^{\gamma^{+}}},\\ a^{S}_{n+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=C_{Q}\frac{(r^{S}_{n}(\tau^{\prime\prime}))^{3+\gamma}l}{2^{n+4}\mathfrak{R}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}a^{S}_{n}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2},\end{cases}$
@@
(5.4) $\begin{split}f(t,x,v)\geq f(t_{1},X_{t_{1},t}(x,v),V_{t_{1},t}(x,v))e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}.\end{split}$
@@
(5.6) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}\\ &\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))ds,\end{split}$
@@
(5.18) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))ds\\ =&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\langle\mathfrak{R}\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\frac{l}{2^{k+3}\mathfrak{R}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v))\\ \geq&a^{S}_{k+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v)).\end{split}$
As a result, the induction hypothesis for $n=k+1$ holds. As a result, (5.1) holds for $0\leq n\leq N$ and we conclude the result.
The following corollary is a direct result (with a bit of modification of the proof) of the previous proposition (Corollary 4.2, [2]):
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$. Fix $\tau^{\prime\prime}>0$ as in Proposition 3.6. Then, given $\Delta^{I}_{T}\in(0,\delta_{T}(\tau^{\prime\prime})]$ with $\tau^{A}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T}]$, for all $0<l\leq\delta_{X}$, there exist $a^{S}(l,\tau^{A},\tau^{\prime\prime},\Delta^{I}_{T})>0$ and $\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})\in\left(0,\frac{\tau^{A}-\tau^{\prime\prime}}{3}\right]$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$(and $L_{f,p}$ if $\gamma<0$) such that we have the following: for $(t,x,v)\in[\tau^{A},\tau^{\prime\prime}+\Delta^{I}_{T}]\times\overline{\Omega}\times\mathbb{R}^{3}$, if for some $t_{1}\in[\tau^{\prime\prime},t-\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})]$ we have $X_{t_{1},t}(x,v)\in\Omega-\Omega_{l}$, then we have
@@
$\geq a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime}))}(v).$
Here, we used the fact that $r^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(\tau^{\prime\prime})\geq 4\mathfrak{R}_{\min}(\tau^{\prime\prime})$ for the last line.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B. Fix $d$ with $0<d<\min\{1,\delta\}$ as in Lemma 2.1. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$ and let $\tau^{\prime\prime}>0$ be as given in Proposition 3.6. Then, there exist $\delta^{B}_{T}(\tau^{\prime\prime})>0$ depending on $\Omega$, $M$, $E_{f}$, and a modulus of continuity of $f$, such that the following holds: For any $\tau^{B}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}^{B}(\tau^{\prime\prime})]$, there exist $a^{B}_{0}(\tau^{B})>0$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$, such that
@@
Then, we recall the notation $t_{\epsilon}(v_{M})$ in (2.23) and define
@@
We next apply the induction assumption and Lemma 3.2 to derive that
@@
By (5.40), it follows that $x_{i}\in\Omega_{\frac{d}{8}}$, so by Remark 3, there exists $1\leq k(i)\leq m_{1}$ such that $x_{i}\in B(x^{0}_{k(i)},\frac{d}{4})$. Notice that $|V_{s,t}(x,v)-v|\geq\frac{\delta_{V}(\tau^{\prime\prime})}{40}$ for some $\tau^{\prime\prime}\leq s\leq t$ Using Lemma 2.6, we deduce that for any $0\leq\tau_{2}\leq t-\tau^{\prime\prime}$, we have
@@
(5.48) $a^{S}_{n,\tau^{\prime\prime}}(\tau^{S}):=\begin{cases}\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{B}),b^{B}_{\tau^{\prime\prime}}(\tau^{B})\right\},\ &n=0\\ (a^{S}_{n-1,\tau^{\prime\prime}}(\tau^{S}))^{2}C_{Q}(r^{\prime\prime}_{n-1,S})^{3+\gamma}\xi_{n}^{\frac{1}{2}}\frac{\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})}{2^{n+1}r^{\prime\prime}_{n,S}}\\ \times e^{-C_{L}\frac{(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left\langle r^{\prime\prime}_{n,S}\right\rangle^{\gamma^{+}}}{2^{n}r^{\prime\prime}_{n,S}}}\ &n\geq 1,\end{cases}$
@@
The proof is the same as the one of Proposition 4.2. We prove by induction on $n$. The base case holds by applying Proposition 5.3: we have
@@
(5.52) $\begin{split}&f(t,x,v)\\ \geq&\int_{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}})}^{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+2}r^{\prime\prime}_{k+1,S}})}e^{-C_{L}(t-s)\langle r^{\prime\prime}_{k+1,S}\rangle^{\gamma^{+}}}\\ &Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))ds.\end{split}$

### v4

@@ 4. Maxwellian bound for the cut-off case for non-fully specular reflection condition
Notice that we used the fact that $\lambda\leq 1$. Hence, by defining $\mathcal{A}_{2}(\tau_{2}):=\lambda^{2}\mathcal{A}_{1}(\tau_{2})$, we conclude the proof of Theorem 1.1 in the case $0\leq\alpha<1$ where
$t\in\left[\tau^{\prime\prime}+\min\left\{\frac{\delta_{T,D}(\tau^{\prime\prime})-\tau^{\prime\prime}}{2},\tau^{\prime\prime}\right\},\tau^{\prime\prime}+\delta_{T,D}(\tau^{\prime\prime})\right].$
To deduce the Theorem 1.2 from Theorem 1.1. We observe that
We consider the first term of (1.12) and (1.13). Then we use Lemma 3.1 to control the damping effect along the characteristic line $X_{s,t}(x,v)$. Thus, the solution is a super solution of the damp transport equation. That is, we either have
(4.47) $\begin{split}f(t,x,v)&\geq f_{0}(X_{0,t}(x,v),v)e^{-tC_{L}(1+|v|^{2})},\end{split}$
when $t\leq t_{\partial}(x,v)$, or
(4.48) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\hskip 14.22636pte^{-t_{\partial}(x,v)C_{L}(1+|v|^{2})}\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\hskip 14.22636pt\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}e^{-t_{\partial}(x,v)C_{L}(1+|v|^{2})},\end{split}$
when $t\geq t_{\partial}(x,v)$. Then Theorem 1.2 can be deduced from the comparison principle.
@@ 5. Maxwellian bound in the cut-off case for fully specular reflection condition
In this chapter, we consider the fully specular reflection condition $\alpha=1$.
@@
(5.3) $\begin{cases}a^{S}_{0}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=a_{0}(\tau^{\prime\prime})e^{-C_{L}(\tau^{\prime}-\tau^{\prime\prime})\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}},\\ a^{S}_{n+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})=C_{Q}\frac{(r^{S}_{n}(\tau^{\prime\prime}))^{3+\gamma}l}{2^{n+4}\mathfrak{R}}e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{n}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2},\end{cases}$
@@
(5.4) $\begin{split}f(t,x,v)\geq f(t_{1},X_{t_{1},t}(x,v),V_{t_{1},t}(x,v))e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}.\end{split}$
@@
(5.6) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\\ &\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds,\end{split}$
@@
(5.18) $\begin{split}&f(t,x,v)\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))\,ds\\ =&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\int_{t_{1}+\frac{l}{2^{k+3}\mathfrak{R}}}^{t_{1}+\frac{l}{2^{k+2}\mathfrak{R}}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{s,t}(x,v))\,ds\\ \geq&e^{-(\tau^{\prime}-\tau^{\prime\prime})C_{L}\left\langle\mathfrak{R}\right\rangle^{\gamma^{+}}}a^{S}_{k}\left(\frac{l}{8},\tau^{\prime\prime},\tau^{\prime},\mathfrak{R}\right)^{2}C_{Q}(r^{S}_{k})^{3+\gamma}\frac{1}{2}\frac{l}{2^{k+3}\mathfrak{R}}\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v))\\ \geq&a^{S}_{k+1}(l,\tau^{\prime\prime},\tau^{\prime},\mathfrak{R})\mathbf{1}_{B(\overline{v}_{i}(\tau^{\prime\prime}),r^{S}_{k+1})}(V_{t_{1},t}(x,v)).\end{split}$
As a result, the induction hypothesis for $n=k+1$ holds. By induction, we deduce that (5.1) holds for $0\leq n\leq N$ and we conclude the result.
The following corollary is obtained by a slight modification of the previous proposition (Corollary 4.2, [2]):
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B with $\nu<0$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$. Fix $\tau^{\prime\prime}>0$ as in Proposition 3.6. Then, given $\Delta^{I}_{T}\in(0,\delta_{T}(\tau^{\prime\prime})]$ with $\tau^{A}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T}]$, for all $0<l\leq\delta_{X}(\tau^{\prime\prime})$, there exist $a^{S}(l,\tau^{A},\tau^{\prime\prime},\Delta^{I}_{T})>0$ and $\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})\in\left(0,\frac{\tau^{A}-\tau^{\prime\prime}}{3}\right]$, depending on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$(and $L_{f,p}$ if $\gamma<0$) such that we have the following: for $(t,x,v)\in[\tau^{A},\tau^{\prime\prime}+\Delta^{I}_{T}]\times\overline{\Omega}\times\mathbb{R}^{3}$, if for some $t_{1}\in[\tau^{\prime\prime},t-\widetilde{t}(l,\tau^{\prime\prime},\tau^{A},\Delta^{I}_{T})]$ we have $X_{t_{1},t}(x,v)\in\Omega-\Omega_{l}$, then we have
@@
$\geq a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B\left(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime})\right)}(V_{t_{1},t}(x,v))$
$=a^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(l,\tau^{\prime\prime},\tau^{\prime\prime}+\Delta^{I}_{T},\mathfrak{R}_{N}^{S}(l,\tau^{\prime\prime},\tau^{A}))\mathbf{1}_{B(0,3\mathfrak{R}_{\min}(\tau^{\prime\prime}))}(v).$
Here, we used the fact that $r^{S}_{\widetilde{N}_{l}(\tau^{\prime\prime},\tau^{A})}(\tau^{\prime\prime})\geq 4\mathfrak{R}_{\min}(\tau^{\prime\prime})\geq 3\mathfrak{R}_{\min}(\tau^{\prime\prime})+|\overline{v}_{i}(\tau^{\prime\prime})|$ for the second line.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the collision kernel $B$ satisfies Assumption B. Fix $d$ with $0<d<\min\{1,\delta\}$ as in Lemma 2.1. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) with $\alpha=1$ and let $\tau^{\prime\prime}>0$ be as given in Proposition 3.6. The following holds: For any $\tau^{B}\in(\tau^{\prime\prime},\tau^{\prime\prime}+\delta_{T}^{B}(\tau^{\prime\prime})]$, there exist $a^{B}_{0}(\tau^{B})>0$, which depends on $\tau^{\prime\prime}$, $\Omega$, $M$, and $E_{f}$, such that
@@
Then, we recall the notation $t_{\epsilon}(v_{M})$ in (2.23) and $\delta_{T}^{B}(\tau^{\prime\prime})$ in (5.32)
@@
Here recall the definition of $\delta_{T}^{B}(\tau^{\prime\prime})$ and $\delta_{T}(\tau^{\prime\prime})$ from (5.32) and (3.61). We next apply the induction assumption and Lemma 3.2 to derive that
@@
By (5.41), it follows that $x_{i}\in\Omega_{\frac{d}{8}}$, so by Remark 2.2, there exists $1\leq k(i)\leq m_{1}$ such that $x_{i}\in B(x^{0}_{k(i)},\frac{d}{4})$. Notice that $|V_{s,t}(x,v)-v|\geq\frac{\delta_{V}(\tau^{\prime\prime})}{40}$ for some $\tau^{\prime\prime}\leq s\leq t$ Using Lemma 2.6, we deduce that for any $0\leq\tau_{2}\leq t-\tau^{\prime\prime}$, we have
@@
(5.49) $a^{S}_{n,\tau^{\prime\prime}}(\tau^{S}):=\begin{cases}\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{S}),b^{B}_{\tau^{\prime\prime}}(\tau^{S})\right\},\ &n=0\\ (a^{S}_{n-1,\tau^{\prime\prime}}(\tau^{S}))^{2}C_{Q}(r^{\prime\prime}_{n-1,S})^{3+\gamma}\xi_{n}^{\frac{1}{2}}\frac{\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})}{2^{n+1}r^{\prime\prime}_{n,S}}\\ \times e^{-C_{L}\frac{(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))\left\langle r^{\prime\prime}_{n,S}\right\rangle^{\gamma^{+}}}{2^{n}r^{\prime\prime}_{n,S}}}\ &n\geq 1,\end{cases}$
@@
The proof is the same as the one of Proposition 4.2. We prove by induction on $n$. The base case holds by applying Proposition 5.3 with the choice $\tau^{B}:=\tau^{S}$: we have
(5.51) $f(t,x,v)\geq\min\left\{a^{B}_{N_{B}(\tau^{\prime\prime}),\tau^{\prime\prime}}(\tau^{S}),b^{B}_{\tau^{\prime\prime}}(\tau^{S})\right\}\mathbf{1}_{B(0,1)}(v)$
@@
(5.53) $\begin{split}&f(t,x,v)\\ \geq&\int_{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+1}r^{\prime\prime}_{k+1,S}})}^{\widetilde{\tau}(\tau^{\prime\prime},\tau^{S})+(\tau^{S}-\widetilde{\tau}(\tau^{\prime\prime},\tau^{S}))(1-\frac{1}{2^{k+2}r^{\prime\prime}_{k+1,S}})}e^{-C_{L}(t-s)\left\langle r^{\prime\prime}_{k+1,S}\right\rangle^{\gamma^{+}}}\\ &Q^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds.\end{split}$